darcy and ian live 24km apart. One day, they agreed to ride a bicycle and meet in a place between the towns where they live. Darcy travelled with a speed of 20kph while Ian 12kph. If they left their respective houses at 9:30, at what time did they meet?
Answer by ankor@dixie-net.com(22740)

darcy and ian live 24km apart.
One day, they agreed to ride a bicycle and meet in a place between the
towns where they live.
Darcy traveled with a speed of 20kph while Ian 12kph.
If they left their respective houses at 9:30, at what time did they meet?
:
Let t = the travel time of the two bikes
:
When they meet, the sum of their travel distances will be 24 km
Write a distance equation; dist = speed * time
:
20t + 12t = 24
32t = 24
t = 24/32 = 3/4 hr which is 45 minutes
:
9:30 + :45 = 10:15 when they meet